Find the Euclidean distance between

(i) (2, 5) and (4, 2)

(ii) (1, 3, 4) and (4, 2, -5)



1
Expert's answer
2021-04-06T07:03:48-0400

"(2,5)-(4,2)=-2,3"


Euclidean d"= \sqrt{(-2)^2+3^2}=\sqrt{12}"




"(1, 3, 4) - (4, 2, -5)=-4,1,9"


Euclidean d ="\sqrt{(-4)^2+1^2+9^2}=\sqrt{98}"